40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 202 of 335  Not logged in ELOG logo
ID Date Author Type Category Subject
  6710   Tue May 29 21:34:08 2012 JenneUpdateSUSETMY oplev spot size reduced

Quote:

Quote:

Quote:

 

 The typical sign of a dying gas laser is that it glows for a few minutes only. The power supplies are fine.

Two new  JDS - Uniphase 1103P lasers ( NT64-104 )  arriving on Monday, May 21

 Yesterday I swapped in new He/Ne laser with output power 3.5 mW  The return spot on qpd is large ~6mm in diameter and 20,500 counts

The spot size reduction require similar layout as ETMX oplev.

 The oplev path is relayed and the spot size on the qpd is reduced. I still have to clean up and replace "Miki Mouse" lens holder.

 Flipped the sign on the ETMY oplev servo gain, since it was wrong.  (It was "-" for both, now it is "+" for both)

  6709   Tue May 29 21:05:30 2012 yutaUpdateIOOPMC, MC alignment are shit

Quote:

  [Keiko, Jenne]

PMC aligned.  Suresh is fixing the measure MC spot positions script, then we'll remeasure MC spot positions.

 [Suresh, Jenne, Yuta]

We measured the MC spot positions twice tonight. Procedure for measuring them is in elog #6688.
The results were;

spot positions in mm (MC1,2,3 pit MC1,2,3 yaw):
    3.3359    3.9595    2.3171   -7.7424   -0.8406    6.4884

spot positions in mm (MC1,2,3 pit MC1,2,3 yaw):
    3.2681    4.0052    2.2808   -7.3965   -0.7624    7.1302

The spot moved by about 0.5 mm since May 25, but we concluded that this displacement is negligible and difficult to be fixed by aligning PSL beam.

We'll align Y arm and X arm next.

  6708   Tue May 29 19:50:01 2012 JenneUpdateIOOPMC, MC alignment are shit

Quote:

The followings are a kind of daily check. Do this without any notice:

- Align PMC.

Quote:

[Yuta, Jenne]

PMC and MC alignment are both shit, although with the WFS on, the MC is pretty good.  We're leaving it for now, so that (a) we don't mess up Koji's work, and (b) we can work on the Xarm.  Steve is doing Yarm oplev stuff, so we'll do Yarm later.

 

 [Keiko, Jenne]

PMC aligned.  Suresh is fixing the measure MC spot positions script, then we'll remeasure MC spot positions.

  6707   Tue May 29 17:40:45 2012 JenneUpdateIOOPMC, MC alignment are shit

Quote:

[Yuta, Jenne]

PMC and MC alignment are both shit, although with the WFS on, the MC is pretty good.  We're leaving it for now, so that (a) we don't mess up Koji's work, and (b) we can work on the Xarm.  Steve is doing Yarm oplev stuff, so we'll do Yarm later.

 [Yuta, Jenne, Suresh]

We pushed on the MC SUS connectors at the back of the rack, and that helped bring MC3 back to where it should be.  Then we looked at MC RFPD DC, and adjusted the optics with the WFS off, so that the refl is ~0.56.  Then when we turn the WFS on, the alignment doesn't really change, so we have offloaded the WFS. 

Now we're measuring the spot positions to check where the MC is.  Then we'll align the arms, and align the green to the arms.

  6706   Tue May 29 15:55:22 2012 steveUpdateSUSETMY oplev spot size reduced

Quote:

Quote:

 

 The typical sign of a dying gas laser is that it glows for a few minutes only. The power supplies are fine.

Two new  JDS - Uniphase 1103P lasers ( NT64-104 )  arriving on Monday, May 21

 Yesterday I swapped in new He/Ne laser with output power 3.5 mW  The return spot on qpd is large ~6mm in diameter and 20,500 counts

The spot size reduction require similar layout as ETMX oplev.

 The oplev path is relayed and the spot size on the qpd is reduced. I still have to clean up and replace "Miki Mouse" lens holder.

There was no IP-ANG coming out of the chamber at this time!

Attachment 1: ETMYoplev.jpg
ETMYoplev.jpg
Attachment 2: IMG_1243.JPG
IMG_1243.JPG
  6705   Tue May 29 15:49:02 2012 KojiUpdateLockingETMX trans camera

Quote:

...will be helpful for acquiring lock after the vent.  We should install a camera at ETMX.

 Do that.

  6704   Tue May 29 15:48:31 2012 KojiUpdateIOOPMC, MC alignment are shit

The followings are a kind of daily check. Do this without any notice:

- Align PMC.

- Check MC spot position with the script (where is it located?). Ignore MC2 result as it can be arbitrary set.

- If the MC1/MC3 spots have moved it means that the PSL beam has moved. If the beam has moved, we should have some discussion what we should do.

- If the spot positions are about the same as before, align the MC mirrors. This should be done by nulling the WFS feedback. (Someone should make a simple script for this WFS offloading.)

------------

Then, start locking both arms

Quote:

[Yuta, Jenne]

PMC and MC alignment are both shit, although with the WFS on, the MC is pretty good.  We're leaving it for now, so that (a) we don't mess up Koji's work, and (b) we can work on the Xarm.  Steve is doing Yarm oplev stuff, so we'll do Yarm later.

 

  6703   Tue May 29 15:29:16 2012 JamieUpdateComputerslatest pynds installed on all new control room machines

The DASWG lscsoft package repositories have a lot of useful analysis software.  It is all maintained for Debian "sqeeze", but it's mostly installable without modification on Ubuntu 10.04 "lucid" (which is based on Debian squeeze).  Basically the only thing that needs to access the lscsoft repositories is to add the following repository file:

controls@rossa:~ 0$ cat /etc/apt/sources.list.d/lscsoft.list 
deb http://www.lsc-group.phys.uwm.edu/daswg/download/software/debian/ squeeze contrib
deb-src http://www.lsc-group.phys.uwm.edu/daswg/download/software/debian/ squeeze contrib

deb http://www.lsc-group.phys.uwm.edu/daswg/download/software/debian/ squeeze-proposed contrib
deb-src http://www.lsc-group.phys.uwm.edu/daswg/download/software/debian/ squeeze-proposed contrib
controls@rossa:~ 0$ 

A simple "apt-get update" then makes all the lscsoft packages available.

lscsoft includes the nds2 client packages (nds2-client-lib) and pynds (python-pynds).  Unfortunately the python-pynds debian squeeze package currently depends on libboost-python1.42, which is not available in Ubuntu lucid.  Fortunately, pynds itself does not require the latest version and can use what's in lucid.  I therefore rebuilt the pynds package on one of the control room machines:

$ apt-get install dpkg-dev devscripts debhelper            # these are packages needed to build a debian/ubuntu package
$ apt-get source python-pynds                              # this downloads the source of the package, and prepares it for a package build
$ cd python-pynds-0.7
$ debuild -uc -us                                          # this actually builds the package
$ ls -al ../python-pynds_0.7-lscsoft1+squeeze1_amd64.deb
-rw-r--r-- 1 controls controls 69210 2012-05-29 11:57 python-pynds_0.7-lscsoft1+squeeze1_amd64.deb

I then copied the package into a common place:

/ligo/apps/debs/python-pynds_0.7-lscsoft1+squeeze1_amd64.deb

I then installed it on all the control room machines as such:

$ sudo apt-get install libboost-python1.40.0 nds2-client-lib python-numpy   # these are the dependencies of python-pynds
$ sudo dpkg -i /ligo/apps/debs/python-pynds_0.7-lscsoft1+squeeze1_amd64.deb

I did this on all the control room machines.

It looks like the next version of pynds won't require us to jump through these extra hoops and should "just work".

  6702   Tue May 29 14:59:39 2012 JenneUpdateIOOPMC, MC alignment are shit

[Yuta, Jenne]

PMC and MC alignment are both shit, although with the WFS on, the MC is pretty good.  We're leaving it for now, so that (a) we don't mess up Koji's work, and (b) we can work on the Xarm.  Steve is doing Yarm oplev stuff, so we'll do Yarm later.

  6701   Tue May 29 14:58:28 2012 JenneUpdateLockingETMX trans camera

...will be helpful for acquiring lock after the vent.  We should install a camera at ETMX.

  6700   Tue May 29 10:08:21 2012 steveUpdateIOOcentered IOO monitoring qpds

IOO Angle & IOO Position QPDs centered.

Attachment 1: 40dtrendIOO.png
40dtrendIOO.png
Attachment 2: IOOcentered.png
IOOcentered.png
  6699   Tue May 29 00:53:57 2012 DenUpdateCDSproblems

I've noticed several CDS problems:

  1. Communication sign on C1SUS model turns to red once in a while. I press diag reset and it is gone. But after some time comes back.
  2. On C1LSC machine red "U" lamp shines with a period ~5 sec.
  3. I was not able to read data from the SR785 using netgpibdata.py. Either connection is not established at all, or data starts to download and then stops in the middle. I've checked the cables, power supplies and everything, still the same thing.
  6698   Tue May 29 00:48:51 2012 DenUpdatePEMsts readout box

STS readout box seems to be partly broken. I've terminated inputs from the seismometer and measured the output. I could not do this for vertical channel because it outputs 7 V DC + 500 mV AC signal. All the switches work fine, 5 V DC is indeed shown when auto zero, calibration, 1 sec resp, sig select are enables. The box has AC power supply that seems to work ok, all measured DC values are equal to the labels. Something is wrong with amplification.

 

sts_box.png

 

  6697   Tue May 29 00:39:52 2012 DenUpdatePEMGuralp noise

I've connected Guralp output to the ADC without readout box. I've got the same noise at low frequencies and even worse noise at high frequencies. However, readout box was still used as DC supply and the signal was read from INPUT test points. I'll do the same experiment without touching readout box at all.

in_out.png

  6696   Tue May 29 00:35:57 2012 DenUpdatePEMguralp readout box

I measured the frequency response of the Guralp readout box and noise by providing sin signal of amplitude 50 mV at 15 Hz for channels 1-3.

1_3.png

It turns out that the gain is ~250, while my liso model simulated it to be 200. This is because it is hard to approximate AD620 amplifier.

Noise of the box does not seem to be too bad at low frequencies.

  6695   Sun May 27 23:15:22 2012 DenUpdatePEMexperiments with seismometers

I wondered if linoleum is a reason of high Guralp noise. I measured Guralp noise in 3 cases: they stand on a very soft piece of paper, linulium and stone.

DSC_4309.JPG               DSC_4310.JPG                  DSC_4311.JPG

I've calibrated the noise to units m/s/sqrt(Hz). Using soft paper we get the worst noise, stone - the best, but noises do not differ that much and still much worse then declared noise in the manual.

loc_noise.png

  6694   Sun May 27 17:19:27 2012 ranaSummaryloreStrawberries

 We have placed some sweet giant strawberries in the fridge; free for eating for anyone working in the lab today or tomorrow:

20120527_145826.jpg

  6693   Sat May 26 23:57:11 2012 DenUpdatePEMGuralp noise

I've looked through the Guralp manual to figure out what noise do they declare. They present it in acceleration units in dB relative to 1 m2 / s4 / Hz. I've converted my measurements to this units and got

noises.png

They declare much better noise. May be linoleum makes an effort. Do we have any isolation boxes?

  6692   Sat May 26 20:40:48 2012 DenUpdatePEMexperiments with seismic noise

I measured relative motion of 2 seismometers separated by 4, 18 and 38 feet.

DSC_4307.JPG         DSC_4308.JPG

Relative motion for difference distances between seismometers is presented below.

    psd_distance.png        ratio.png

 

Al frequencies f < fcrit relative motion becomes smaller then motion of each point. At these frequencies, the ratio relative / absolute motion can be approximated as ratio = C fa . The following table summarizes fcrit, C and a for different length between seismometers.

 

L, ft fcrit, Hz
C a
4 45 0.0074 1.30
18 30 0.0057 1.52
38 15 0.0208 1.43

Using this approximation we can estimate the desired noise floor of the seismometer to subtract seismic noise from MC_F

sim_delta.png

Desired seismometer noise should be 100 times less then Guralp's. ADC noise is still less then this level, so this will not be a problem.

Note: for longer cavities condition for seismometer noise becomes more week as fcrit decreases with length.

  6691   Sat May 26 15:59:19 2012 DenUpdateIOOGuralp noise is high

As I've mentioned in yesterday's elog MC mirrors start to move in a synchronistic manner. I've plotted DELTA_GUR = GUR1_X - alpha * GUR2_X, where alpha = const to make the transfer functions SEISMOMETER -> ADC equal for each channel. I've noticed that DELTA_GUR decreases below 10 Hz compared to GUR1_X as theoretically predicted. But starting from 1 Hz DELTA_GUR starts to increase. I decided that this is Guralp noise floor. Today I checked this, this is indeed the case.

In the frequency range 0.01 - 1.5 Hz Gur noise is comparable to the signal DELTA_GUR. For that reason we see low coherence between MC_F and GUR1_X in this frequency range. 

gur_noise.png    MCF_GUR.png

Guralp noise floor was determined by placing 2 seimometers close to each other and subtracting by Wiener filtering.

DSC_4306.JPG

Conclusion: To filter seismic noise out of MC_F we need more sensitive seimometeres.

 

  6690   Sat May 26 01:28:58 2012 KojiUpdateGeneralClipping improvement

The beam at the dark port was largely clipped.
The arm transmissions (TRX/TRY) were ~0.6 no matter how you try to align.

The PZT1/2 X was scanned and TRX/TRY are now both ~0.9.
This means that the beam was mostly clipped in or after the mode mathing optics.

FPMI contrast is (2.45-0.02)/(2.45+0.02) = 98%
Not so bad although it does not tell anything about clipping before the IFO.

In the next opportunity, we should try to "pin" the beam axis on the arm mirrors using A2L.

1. I am not sure exactly where the clipping was. It means that a small clipping might still be there.

2. I am afraid that the beam is still clipped in the faraday. The beam can be scanned using the steering on the PSL table.

 

Attachment 1: Untitled.png
Untitled.png
  6689   Sat May 26 00:08:41 2012 DenUpdateIOOMC_F low frequency noise

 MC_F low frequency noise might be due to local damping electronics. I did not measure OSEM noise, but even without it electronics (AA -> ICS 110 -> ADC) provide sufficient amount of noise. 

These 2 image show electronics noise and coherence between OSEM signal / seismic

osem_noise.png           gur_suspos_coh.png

From these 2 plots we might think that SNR > 10 and coherence OSEM / GUR is high at the frequency range 0.1 - 10 Hz and this is not a big problem.

However, at low frequencies the length of seismic waves becomes large enough and relative oscillations of MC2 and MC13 decrease.

For 1 wave ( u(MC2) - u(MC1) ) / u(MC2) = sin(2 * pi * L  * f  / c), L - distance between MC1 and MC2 where 2 seismometers are located. So MC123 move according to seismic motion and electronics noise is not seen unless we look at MC Length. Here this noise is seen, because mirrors move in a synchronistic manner. 

To check this I measured seismic noise with 2 guralps at distance 12 meters - at MC1 and MC2. Then I've computed the difference between these signals. And indeed at low frequencies, relative motion is much less.

Green, blue - GUR1,2_X

Red - differential motion GUR1_X - GUR2_X

dgur.png

 

The following plot illustrates how electronics noise effects MC_F. Green is the signal to coils. Red - electronics noise. Blue, black, cyan - simulated contribution to MC_F for different seismic waves speed. Most probably seismic waves have waves in the range 50 - 800 m/s, others are deep. The plot shows that electronics noise is big enough to disturb coherence between MC_F and seismic noise. 

mc_noises.png

Here is a rough calculation of the seismic waves speed. The following plot shows the ratio of psd of differential MC2-MC1 motion to MC2 motion.

ratio.png

If seismometers would be very far, ratio would be 1 if we neglect the difference in transfer function SEISMOMETER -> ADC for each channel. The drift of the ratio from 1 to 1.3 demonstrates this effect. Ratio starts to decrease at 15 Hz according to sin (2*pi*L*f/c) ~ 2*pi*L/c * f. So 2*pi*L/c * f_0 = pi/2 => c = 4 * L * f = 600 m / sec.

  6688   Fri May 25 23:11:50 2012 SureshUpdateIOOMC spot positions measured

[Koji, Yuta, Suresh]

We measured the MC spot positions after re-aligning the MC.  The spot positions are listed below:

spot positions in mm (MC1,2,3 pit MC1,2,3 yaw):
    3.9073    6.6754    2.8591   -7.6985   -0.9492    7.0423

 

Procedure:

1) In the directory /opt/rtcds/caltech/c1/scripts/ASS/MC  we have the following scripts

      a) mcassUp:    This sets up the MCASS lockins to excite each of the MC mirrors at a different frequency

      b) mcassOn:    This sets the MCASS output matrix to actually send the excitation signals to the mirrors

      c) senseMCdecenter:  This sequentially introduces a 10% offset into the coil gains of each mirror degree of freedom.   It also sends the lockin output data to the screen. 

      d) sensemcass.m :  This is a matlab file which digests the data gathered by the senseMCdecenter script to print a couple of plots and compute the spot positions.

      e) MCASS_StripTool.stp:  This is a set-up file for the StripTool which allows us to see the MCASS-lockin_outputs.  It is nice to see the action of senseMCdecenter  script  at work.

 

2) So the series of commands to use are

      a) ./StripTool  <-- MCASS_StripTool.stp

      b) ./mcassUp

      c) ./mcassOn

      d) ./senseMCdecenter | tee Output_file

       e) ./mcassOff

      f) ./mcassDown

       g) matlab <-- sensemcass.m  <---- Output_file

 

  6687   Fri May 25 20:45:25 2012 Duncan MacleodSummaryComputer Scripts / Programs40m summary webpages

There is now a job in the crontab that will run the shell wrapper every hour, so the pages _should_ take care of themselves. If you make adjustments to the configuration file they will get picked up on the hour, or you can just run the script by hand at any time.

$ crontab -l
# m h  dom mon dow   command
0 */1 * * * bash /home/controls/public_html/summary/bin/c1_summary_page.sh > /dev/null 2>&1

  6686   Fri May 25 19:13:10 2012 Duncan MacleodSummaryComputer Scripts / Programs40m summary webpages

40m summary webpages

 The aLIGO-style summary webpages are now running on 40m data! They are running on megatron so can be viewed from within the martian network at:

http://192.168.113.209/~controls/summary

At the moment I have configured the 5 seismic BLRMS bands, and a random set of PSL channels taken from a strip tool.

Technical notes

  • The code is in python depending heavily on the LSCSoft PyLAL and GLUE modules.
    • /home/controls/public_html/summary/bin/summary_page.py
  • The HTML is supported by a CSS script and a JS script which are held locally in the run directory, and JQuery linked from the google repo.
    • /home/controls/public_html/summary/summary_page.css
    • /home/controls/public_html/summary/pylaldq.js
  • The configuration is controlled via a single INI format file
    • /home/controls/public_html/summary/share/c1_summary_page.ini

Getting frames

Since there are no segments or triggers for C1, the only data sources are GWF frames. These are mounted from the framebuilder under /frames on megatron. There is a python script that takes in a pair of GPS times and a frame type that will locate the frames for you. This is how you use it to find T type frames (second trends) for May 25 2012:

python /home/controls/public_html/summary/bin/framecache.py --ifo C1 --gps-start-time 1021939215 --gps-end-time 1022025615 --type T -o framecache.lcf

If you don't have GPS times, you can use the tconvert tool to generate them

$ tconvert May 25
1021939215

The available frame types, as far as I'm aware are R (raw), T (seconds trends), and M (minute trends).

Running the code

The code is designed to be fairly easy to use, with most of the options set in the ini file. The code has three modes - day, month, or GPS start-stop pair. The month mode is a little sketchy so don't expect too much from it. To run in day mode:

python /home/controls/public_html/summary/bin/summary_page.py --ifo C1 --config-file /home/controls/public_html/summary/share/c1_summary_page.ini --output-dir . --verbose --data-cache framecache.lcf -SRQDUTAZBVCXH --day 20120525

Please forgive the large apparently arbitrary collection of letters, since the 40m doesn't use segments or triggers, these options disable processing of these elements, and there are quite a few of them. They correspond to --skip-something options in long form. To see all the options, run

python /home/controls/public_html/summary/bin/summary_page.py --help

There is also a convenient shell script that will run over today's data in day mode, doing everything for you. This will run framecache.py to find the frames, then run summary_page.py to generate the results in the correct output directory. To use this, run

bash /home/controls/public_html/summary/bin/c1_summary_page.sh

Configuration

Different data tabs are disabled via command link --skip-this-tab style options, but the content of tabs is controlled via the ini file. I'll try to give an overview of how to use these. The only configuration required for the Seismic BLRMS 0.1-0.3 Hz tab is the following section:

 

[data-Seismic 0.1-0.3 Hz]
channels = C1:PEM-RMS_STS1X_0p1_0p3,C1:PEM-RMS_STS1Y_0p1_0p3,C1:PEM-RMS_STS1Z_0p1_0p3
labels = STS1X,STS1Y,STS1Z
frame-type = R
plot-dataplot1 =
plot-dataplot3 =
amplitude-log = True
amplitude-lim = 1,500
amplitude-label = BLRMS motion ($\mu$m/s)

The entries can be explained as follows:

  1. '[data-Seismic 0.1-0.3 Hz] - This is the section heading. The 'data-' mark identifies this as data, and is a relic of how the code is written, the 'Seismic 0.1-0.3 Hz' part is the name of the tab to be displayed in the output.
  2. 'channels = ...' - This is a comma-separated list of channels as they are named in the frames. These must be exact so the code knows how to find them.
  3. 'labels = STS1X,STS1Y,STS1Z' - This is a comma-separated list of labels mapping channel names to something more readable for the plots, this is optional.
  4. 'frame-type = R' - This tells the code what frame type the channels are, so it can determine from which frames to read them, this is not optional, I think.
  5. 'plot-dataplotX' - This tells the code I want to run dataplotX for this tab. Each 'dataplot' is defined in it's own section, and if none of these options are given, the code tries to use all of them. In this configuration 'plot-dataplot1' tells the code I want to display the time-series of data for this tab.
  6. 'amplitude-XXX = YYY' - This gives the plotter specific information about this tab that overrides the defaults defined in the dataplotX section. The options in this example tell the plotter that when plotting amplitude on any plot, that axis should be log-scale, with a limit of 1-500 and with a specific label. The possible plotting configurations for this style of option are: 'lim', 'log', 'label', I think.

Other compatible options not used in this example are:

 

  • scale = X,Y,Z - a comma-separated list of scale factors to apply to the data. This can either be a single entry for all channels, or one per channel, nothing in between.
  • offset = X,Y,Z - another comma-separate list of DC offsets to apply to the data (before scaling, by default). DAQ noise may mean a channel that should read zero during quick times is offset by some fixed amount, so you can correct that here. Again either one for all channels, or one per channel.
  • transform = lambda x: f(x) - a python format lambda function. This is basically any mathematical function that can be applied to each data sample. By default the code constructs the function 'lambda d: scale * (d-offset)', i.e. it calibrates the data by removing the offset an applying the scale.
  • band = fmin, fmax - a low,high pair of frequencies within which to bandpass the data. Sketchy at best...
  • ripple_db = X - the ripple in the stopband of the bandpass filter
  • width = X - the width in the passband of the bandpass filter
  • rms_average = X - number of seconds in a single RMS average (combine with band to make BLRMS)
  • spectrum-segment-length = X - the length of FFT to use when calculating the spectrum, as a number of samples
  • spectrum-overlap = X - the overlap (samples) between neighbouring FFTs when calculating the spectrum
  • spectrum-time-step = X - the length (seconds) of a single median-mean average for the spectrogram

At the moment a package version issue means the spectrogram doesn't work, but the spectrum should. At the time of writing, to use the spectrum simple add 'plot-dataplot2'.

You can view the configuration file within the webpage via the 'About' link off any page.

Please e-mail any suggestions/complaints/praise to duncan.macleod@ligo.org.

  6685   Fri May 25 17:52:08 2012 JamieUpdateComputersallegra now running Ubuntu 10.04

The last of the control room machines is now upgraded.

  6684   Fri May 25 17:50:38 2012 JamieUpdateComputersASS scripts on new ubuntu machines

Quote:

[Den, Yuta]

Background:
 ASS and many other scripts don't work on new ubuntu machines.

What we did:
1. Installed C-shell on rossa and rosalba(Ubuntu machine).
  sudo apt-get insall csh

2. Find out that
  /opt/rtcds/caltech/c1/scripts/AutoDither/alignY

runs, but
  /opt/rtcds/caltech/c1/scripts/medmrun /opt/rtcds/caltech/c1/scripts/AutoDither/alignY

doesn't run. It gives us the following error messages.

ezcawrite: error while loading shared libraries: libca.so: cannot open shared object file: No such file or directory
ezcaswitch: error while loading shared libraries: libca.so: cannot open shared object file: No such file or directory

Result:
 ASS scripts run on rossa and rosalba, but not with medmrun.
 At least ASS scripts run on pianosa(ubuntu machine) with medmrun. So we decided to wait for JAMIE to fix it.

Apparently the environment was not being properly inherited by the scripts launched from medmrun.  We modified the medmrum script so that it executes things with an interactive shell ("bash -i -c ...") and this fixed the problem (by assuring that it sources all the interactive environment configs (i.e. ~/.bashrc)).  I'm still not sure why we were seeing different behavior on pianosa, but at least the solution we have now should be robust.

As a reminder, all scripts launched from MEDM should use medmrun:

/opt/rtcds/caltech/c1/scripts/medmrun
  6683   Fri May 25 16:58:54 2012 JamieConfigurationComputers.bashrc for workstations

I have setup a shared .bashrc for all the workstations that is symlinked to the normal location on all machines:

controls@rossa:~ 0$ ls -al /home/controls/.bashrc 
lrwxrwxrwx 1 controls controls 23 2012-05-25 15:37 /home/controls/.bashrc -> /users/controls/.bashrc
controls@rossa:~ 0$ 

This should help simplify maintenance considerably.  Editing that file on one machine will edit it for all.  Just edit this one file!  Don't try to get fancy and add extra files!

I also added a bunch of aliases that had previously been missing.  This should help with some of the problems that people had been having.

NOTE: PLEASE DO NOT CHANGE THE DEFAULT SHELL!  We are using bash, because that's what the sites are now using and we want to be as compatible as possible.

You can of course still write scripts in csh/tcsh or use tcsh in a shell if you wish.   Just don't change the default shell for the controls user.

  6682   Fri May 25 16:10:53 2012 KojiUpdateGeneralStripTool on Zita restored

ssh -X pianosa

On pianosa

StripTool PSL.strip&
StripTool IOO.strip&
StripTool seismic.strip&

 

  6681   Fri May 25 13:24:48 2012 DenUpdateSUSMC3

AA IN -> COIL DRIVER IN transfer function for MC3

freq_resp.png

I've provided excitation to the AA input, the same for all OSEM channels. In the digital domain coherence between C1:SUS-MC3_ULSEN_INMON / C1:SUS-MC3_ULCOIL_INMON and other channels OSEM -> COIL is 1 starting from 0.1Hz.

test_dig.png

The only thing left to understand is why the coherence AA IN / COIL DRIVER IN measured in the analog domain is not 1 in the frequency range 0.1 - 1 Hz. It does not look like just SRS noise. I've connected Ch 1 and 2 to the source, coherence is close to 1.

  6680   Fri May 25 02:56:44 2012 DenUpdateSUSMC3 local damping

I've terminated input to AA filters and measured signal to coils C1:SUS-MC3_??COIL_OUT.

DSC_4305.JPG

 

I compared this noise to the signal when OSEM are connected to ADC.

    osem_noise.png

 

I made BNC -> LEMO board such that all LEMOs have the same signal equal to BNC signal. I provided excitation of 50 mV as white noise to the input of the AA filter and measured coherence between excitation and MC3 coil driver. The path is

AA -> ICS 110 -> Pentium -> Pentek -> AI -> Univ Dewhitening -> Coil Driver

As all inputs have the same signal, matrices that recombine the signals should not affect coherence. But what I got for coherence between AA IN / Dewhitening OUT is

osem_coil.png

  6679   Thu May 24 19:39:18 2012 SureshUpdateIOOMC and WFS alignment adjusted

[Yuta, Suresh]

We found that the MC was not locking and that the alignment between PSL and MC was too poor to obtain a TEM00 mode in the MC.   To correct the situation we went through the following steps:

1) We burt restored the MC alignment slider values to their values at 3:07 AM of today

2) We turned off the MC-autolocker and the ASC signal to the coils.   Then aligned the PSL beam into the MC (with the MC servo loop off) to obtain the TEM00 mode.  We had to adjust the zig-zag at the PSL output by quite a bit to maximise MC transmission.

3) We then centered the spot on the MC2 face and centered the transmitted beam on the MC2_TRANS_QPD

4) Next, we centered the beams on the MC_WFS sensors.

5) Turning on the WFS loops after this showed that everything works fine and WFS loops do not accumulate large offsets.

 

 

  6678   Thu May 24 16:39:47 2012 steveUpdateGeneralsteve's webpage example

3 strip charts monitoring on 24 hours time scale: Vac,  PSL-IFO,  SUS and 10 channel video monitoring inside - outside  of 40m lab

Attachment 1: 05241201.PDF
05241201.PDF
  6677   Thu May 24 16:13:05 2012 yutaUpdateComputersASS scripts on new ubuntu machines

[Den, Yuta]

Background:
 ASS and many other scripts don't work on new ubuntu machines.

What we did:
1. Installed C-shell on rossa and rosalba(Ubuntu machine).
  sudo apt-get insall csh

2. Find out that
  /opt/rtcds/caltech/c1/scripts/AutoDither/alignY

runs, but
  /opt/rtcds/caltech/c1/scripts/medmrun /opt/rtcds/caltech/c1/scripts/AutoDither/alignY

doesn't run. It gives us the following error messages.

ezcawrite: error while loading shared libraries: libca.so: cannot open shared object file: No such file or directory
ezcaswitch: error while loading shared libraries: libca.so: cannot open shared object file: No such file or directory

Result:
 ASS scripts run on rossa and rosalba, but not with medmrun.
 At least ASS scripts run on pianosa(ubuntu machine) with medmrun. So we decided to wait for JAMIE to fix it.

  6676   Thu May 24 15:10:43 2012 SureshSummaryGeneralIOO (MC) health check webpage layout

   Here is the suggested layout of the MC health check web page layout. I will update the Omnigraffle file as people comment and suggest changes.  If you want the file let me know.

 

MC_Health_Check.pdf

  6675   Thu May 24 14:49:59 2012 KojiSummaryGeneralDaily news idea

Top tab categolies:

  • Summary
  • CDS
    • CDS Status
  • PEM
    • Seismic 24h trend
    • Accoustic 24h trend
    • Weather/Temp/Barometer/etc 24h trend
  • PSL/IOO
    • PSL summary trend / duty ratio
    • IOO summary (MC Health Check/IOO QPD trends / IFO QPD trends / Transmon QPD trends) duty ratio
  • SUS
    • Summary
    • OSEM PSD/trend
    • OPLEV PSD/trend
  • IFO
    • DC Mon
    • RF ports
    • OMC
  • Steve
    • Vacuum
  • Misc.

IFO

  • DC Monitors
    • Incident beam power trend (24h)
    • AS/REFL/POP/TRX/TRY bean power trend (24h)
    • AS/POP RF beam power trend (24h)
  • RF port
    • DARM sensitivity PSD (mean/min/max/reference) for an hour
    • DARM/CARM/PRCL/MICH/SRCL PSD
    • DARM/CARM/PRCL/MICH/SRCL (freq vs Gaussianity)
    • DARM/CARM/PRCL/MICH/SRCL calibration trend
  • OMC
    • TBD

 

  6674   Thu May 24 13:28:38 2012 DenUpdatePEMHEPA

HEPA filter was running at 90% of max. I reduced it to 20%. Acoustic noise moved down

psl_acoustic.png

The range of MCL oscillations has also decreased but fluctuations in the frequency range 10-100 are still present.

MCL is much more stable now.

mcl_90.png        mcl_20.png

 

  6673   Thu May 24 12:23:00 2012 DenUpdateGeneralconnection error monitor for the website

The main page of connection error monitor can be a scheme of computers and models with connections. If there is an error in connection, the corresponding arrows turns from green to red.

errors_overview.png

 

Each arrow is a line to a more detailed information about the channels.

errors.png

  6672   Thu May 24 10:10:04 2012 steveUpdateIOOPMC locked

 The PMC  behavior is not changed.

 

Attachment 1: pmc4d.png
pmc4d.png
  6671   Thu May 24 02:55:36 2012 DenUpdatePEMacoustic noise in MCL

Mic in the PSL showed that fluctuations in the MCL in the frequency range 10 - 100 Hz are due to acoustic noise. I've measured MCL, MCL / PSL mic coherence 2 times with interval 300 seconds.

Surprisingly, acoustic noise level did not change but MC sometimes is more sensitive to acoustic noise, sometimes less.

audio_noise.png

  6670   Thu May 24 01:17:13 2012 DenUpdateCDSPMC autolocker

Quote:

 

  • SCRIPT
    • Auto-locker for PMC, PSL things - DEN

 I wrote auto-locker for PMC. It is called autolocker_pmc, located in the scripts directory, svn commited. I connected it to the channel C1:PSL-PMC_LOCK.  It is currently running on rosalba. MC autolocker runs on op340m, but I could not execute the script on that machine

op340m:scripts>./autolock_pmc
./autolock_pmc: Stale NFS file handle.

I did several tests, usually, the script locks PMC is a few seconds. However, if PMC DC output has been drift significantly, if might take longer as discussed below.

The algorithm:

       if autolocker if enabled, monitor PSL-PMC_PMCTRANSPD channel
       if TRANS is less then 0.4, start locking:
               disengage PMC servo by enabling PMC TEST 1
               change PSL-PMC_RAMP unless TRANS is higher then 0.4 (*)
               engage PMC servo by disabling PMC TEST 1
       else sleep for 1 sec
 

(*) is tricky. If RAMP (DC offset) is specified then TRANS will be oscillating in the range ( TRANS_MIN, TRANS_MAX ). We are interested only in the TRANS_MAX. To make sure, we estimate it right, TRANS channel is read 10 times and the maximum value is chosen. This works good.

Next problem is to find the proper range and step to vary DC offset RAMP. Of coarse, we can choose the maximum range (-7, 0) and minimum step 0.0001, but it will take too long to find the proper DC offset. For that reason autolocker tries to find a resonance close to the previous DC offset in the range (RAMP_OLD - delta, RAMP_OLD + delta), initial delta is 0.03 and step is 0.003. It resonance is not found in this region, then delta is multiplied by a factor of 2 and so on. During this process RAMP range is controlled not to be wider then (-7, 0).

The might be a better way to do this. For example, use the gradient descent algorithm and control the step adaptively. I'll do that if this realization will be too slow.

I've disabled autolocker_pmc for the night.

  6669   Wed May 23 21:32:15 2012 SureshUpdateIOOWFS didn't turn off automatically

Quote:

I just sat down in the control room, and discovered the PMC (and everything else) unlocked.  I relocked the PMC, but the MC wasn't coming back.  After a moment of looking around, I discovered that the WFS were on, and railing.  I ran the "turn WFS off" script, and the MC came back right away, and the WFS came on as they should.

We need to relook at the WFS script, or the MC down script, to make sure that any time the MC is unlocked, no matter why it unlocked, the WFS output is off and the filter histories are cleared.

    The only script that can currently take this action is the MC autolocker.  If that is disabled first and the PMC unlocks later, the WFS will not be turned off.  During the last round of discussions we had about the autolocker script, sometime last Nov, we decided that too much automation is not desirable and that the autolocker should be kept as simple as possible.

 

  6668   Wed May 23 20:41:37 2012 SureshBureaucracyGeneral40m Meeting Action Items: Tip-tilts : cabling and electronics

Quote:
..........
  • Tip Tilts
    • Prepare electronics for TTs (coil drivers) - JAMIE
    • In-air TT testing to confirm we can control / move TTs before we vent - SURESH
    • Connect TTs to digital system and controls, lay cables if needed - JAMIE with SURESH
  • .....

 

 [Koji, Suresh]

  We tried to locate the sixteen analog output channels we need to control the four tip-tilts (four coils on each).  We have only 8 available channels on the C1SUS machine. 

  So we will have to plug-in a new DAC output card on one of the machines and it would be logical to do that on the C1IOO machine as the active tip-tilts are conceptually part of the IOO sub-system.   We have to procure this card if we do not already have it.  We have to make an interface between this card output and a front panel on the 1X2 rack.  We may have to move some of the sub-racks on the 1X2 rack to accommodate this front panel.

  We checked out the availability of cards (De-whitening, Anti-imaging, SOS coil drivers)  yesterday.  In summary: we have all the cards we need (and some spares too).  As the De-whitening and Anti-imaging cards each have 8 channels, we need only two of each to address the sixteen channels.  And we need four of the SOS coil drivers, one for each tip-tilt.  There are 9 slots available on the C1IOO satellite expansion chassis (1X1 rack), where these eight cards could be accommodated.

  There are two 25 pin feed-thoughs, where the PZT drive signals currently enter the BS chamber.  We will have to route the SOS coil driver outputs to these two feed-throughs. 

   Inside the BS chamber, there are cables which carry the PZT signals from the chamber wall to the the table top, where they are anchored to a post (L- bracket).  We need a 25-pin-to-25-pin cable (~2m length) to go from the post to the tip-tilt (one for each tip-tilt).  And then, of course, we need quadrapus cables (requested from Rich) which fit inside each tip-tilt to go to the BOSEMs.

   I am summarising it all here to give an overview of the work involved.

 

 

  6667   Wed May 23 13:57:11 2012 JenneBureaucracyGeneral40m Meeting Action Items

PRIORITY ITEMS:

Align Ygreen beam - JENNE, YUTA

Arm cavity sweeps, mode scan - JENNE, YUTA

ASS doesn't run on Ubuntu! or CentOS Fix it! - YUTA, JENNE, JAMIE's help

Input matricies, output filters to tune SUS.  check after upgrade. - JENNE

POX11 whitening is not toggling the analog whitening??? - JAMIE, JENNE, KOJI

Decide on plots for 40m Summary page - DEN, STEVE, JENNE, KOJI, JAMIE, YUTA, SURESH, RANA, DUNCAN from Cardiff/AEI

 Look into PMC PZT drift - PZT failing? Real MC length change? - JENNE, KOJI, YUTA

THE FULL LIST:

  • General
    • Revamp control room - more monitors, maybe move TV shelves. - JENNE
    • Audio system for the signals!!!! Even a crappy one! - LEO
  • SUS
    • Input matricies, output filters to tune SUS.  check after upgrade. - JENNE
  • IFO
    • Fix occasional common-mode power transient in the arm transmissions. Probably an alignment thing. Would ISS help? - MIKEJ?
    • Drift of the green incident axis -> Assess the amount of the drift / replace the mount - JENNE, KOJI
    • Calibration of POP22 / AS110 - JENNE
    • PMC/IMC/ARM characterization (loss, finesse, reflectivity, etc) - JENNE
    • Align green beam - JENNE, YUTA
    • Arm cavity sweeps, mode scan - JENNE, YUTA
    • Align AS OSA (others?) - JENNE  
    • Investigate PRMI glitches, instability - JENNE
    • PZT or Picomotor mounts for PSL/ALS beams - WHO??
    • ALS on the both arm simultaneously / common / diff ALS scripts - JENNE
  • CDS
    • Capture OSA signals in CDS (the 'scope TDS1001B has a USB port in the back for connecting to the computer) - WHO??
    • Transmon (arms) for high and low power - WHO??
    • POX11 whitening is not toggling the analog whitening??? - JAMIE, JENNE, KOJI
    • Install guardians to monitor EPICS values - WHO??
  • Electronics
    • Actuator noise level characterization (coil driver response in V/m & coil driver noise level V/rtHz) - WHO??
    • Beat box - KOJI
    • I/Q DFD - WHO??
    • Improvement of POP22/110/AS110 RF circuits. - WHO??
  • MEDM
    • Complete 40m overview screen - everything should be clickable with pseudo 3D icons - JENNE
    • Better suspension screen - copy from sites. - WHO??
    • Script to generate a MEDM tree - WHO??
    • Resurrect MEDM snapshots - WHO??
    • New ! buttons on every screen, include wiki page - WHO??
    • Screen for errors in signal transmission between models - DEN
  • SCRIPT
    • Locking scripts with integrator triggers / disabling when unlocked - JENNE
    • Daily diagnosis of the MC spot positions (there must be something already...) - SURESH?
    • Daily/occasional adjustment of the incident axis on the MC - SURESH?
    • IFO_CONFIGURE scripts still do a burt restore, so step the optics.  Need to remove optic alignment from config .snap files, use reg restore script instead. - JENNE
    • OPLEV/OSEM trending script before the IFO work for diagnosis. Put into 40m summary screen. - JENNE, DUNCAN from Cardiff/AEI
    • Auto-locker for arms - JENNE
    • Auto-locker for PMC, PSL things - DEN
    • Diagnostic script for CDS - mx_stream, other stuff. - WHO??
  • Video
    • If each video screen has a caption, that would be great - WHO??
    • GUI interface of "videoswitch" - MIKEJ
    • Mouse Jiggler for Zita (called Swinput?) - JENNE
  • Ubuntu vs. CentOS
    • burttoday is not aliased in ubuntu.    burttoday:      aliased to cd /opt/rtcds/caltech/c1/burt/autoburt/today/ - JAMIE
    • Upgrade Allegra, make sure connect to DTT, Dataviewer, AWG. - JAMIE
    • ASS doesn't run on Ubuntu! or CentOS Fix it! - JENNE, JAMIE's help
    • CentOS machines cannot open simulink models properly (get "Bad Link"s everywhere, so you can't do anything useful) - JAMIE
  • MM
    • IPPOS beam measurement - SURESH with JENNE
    • AS beam measurement (if beam is bright enough) - SURESH and JENNE
    • Mode matching calculations, sensitivity to MC waist measurement errors, PRM position - JENNE
    • Think up diagnostic measurement to determine mode matching to PRC while chambers are open, while we tweak MMT - JAMIE, JENNE, KOJI, SURESH
    • Use sensoray to capture, measure beam mode at AS, POP - YUTA
  • Stray Light
    • Scattered light measurement at the end stations: design / confirmation of the mechanical parts/optics/cameras - JAN
    • Align AUX laser into dark port - KOJI, YUTA
  • OAF
    • OAF comparison plot, both online and offline, comparing static, adaptive and static+adaptive - DEN
    • Static-only OAF noise budget (Adaptive noise budget as next step) - DEN
    • Script for daily / weekly re-calculation of Wiener, post to elog if need changing - DEN
  • Tip Tilts
    • Prepare electronics for TTs (coil drivers) - JAMIE
    • In-air TT testing to confirm we can control / move TTs before we vent - SURESH
    • Connect TTs to digital system and controls, lay cables if needed - JAMIE with SURESH
  • RF Photodiodes
    • Opto Energy diode laser - purchase - ERICG
    • Purchase fibers, splitters, other supplies - ERICG
    • Set everything up - ERICG

 

  6666   Wed May 23 11:30:39 2012 steveUpdatePEMnew emergency back up lights installed

New self checking emergency back up lights were installed at 13 locations in the 40m lab.

Attachment 1: IMG_1217.JPG
IMG_1217.JPG
  6665   Wed May 23 10:40:21 2012 steveUpdateIOOPMC locked

Quote:

I locked the PMC and the MC followed instantly.

 

Attachment 1: pmc.png
pmc.png
  6664   Tue May 22 23:36:02 2012 DenUpdatePEMmircophones

I've put 5 EM172 microphones close together and measured there signal and coherence. They are plugged in to accelerometer channels.

close_micro.png

 

Then I've suspended microphones around the MC - 2 at MC2, 2 at MC1,3 and 1 at PSL. The amplifier box is above STS readout box.

 DSC_4288.JPG           DSC_4299.JPG       DSC_4304.JPG      DSC_4302.JPG

Microphone close to PSL gave a strong coherence with MC_F, as we already saw it using Blue Bird Microphone.

ACC_MC2_XY channels <=> MC2 microphones

ACC_MC1_XY channels <=> MC1,3 microphones

ACC_MC1_Z channel <=> PSL microphones

5_mic.png

 

  6663   Tue May 22 20:46:38 2012 yutaUpdateCDSMEDM suspension screens using macro

I fixed the problem Jamie pointed out in elog #6657 and #6659.

What I did:
1. Created the following template files in /opt/rtcds/userapps/trunk/sus/c1/medm/templates/ directry.

SUS_SINGLE_LOCKIN1.adl
SUS_SINGLE_LOCKIN2.adl
SUS_SINGLE_LOCKIN_INMTRX.adl
SUS_SINGLE_OPTLEV_SERVO.adl
SUS_SINGLE_PITCH.adl
SUS_SINGLE_POSITION.adl
SUS_SINGLE_SUSSIDE.adl
SUS_SINGLE_TO_COIL_MASTER.adl
SUS_SINGLE_COIL.adl
SUS_SINGLE_YAW.adl
SUS_SINGLE_INMATRIX_MASTER.adl
SUS_SINGLE_INPUT.adl
SUS_SINGLE_TO_COIL_X_X.adl
SUS_SINGLE_OPTLEV_IN.adl
SUS_SINGLE_OLMATRIX_MASTER.adl

To open these files, you have to define $(OPTIC) and $(DCU_ID).
For SUS_SINGLE_TO_COIL_X_X.adl, you also have to define $(FILTER_NUMBER), too. See SUS_SINGLE_TO_COIL_MASTER.adl.

2. Fixed the following screens so that they open SUS_SINGLE.adl.

C1SUS_WATCHDOGS.adl
C1IOO_MC_ALIGN.adl
C1IOO_WFS_MASTER.adl
C1IFO_ALIGN.adl

  6662   Tue May 22 20:24:06 2012 JamieUpdateComputersrossa is now running Ubuntu 10.04

Now same as pianosa and rosalba.  I'll upgrade allegra on Friday.

  6661   Tue May 22 20:01:26 2012 DenUpdateCDSerror monitor

I've created transmission error monitors in rfm, oaf, sus, lsc, scx, scy and ioo models. I tried to get data from every channel transmitted through PCIE and RFM. I also included some shared memory channels.

The medm screen is in the EF STATUS -> TEM. It shows 16384 for the channels that come from simulation plant. Others are 0, that's fine.

ELOG V3.1.3-